# Diff-SSL-G-COMP: Towards a large-scale and diverse dataset for virtual analog modeling

### v1

@@ 4.2 Black-Box Methods
Table 2 illustrates the benchmarking results on black-box methods. It can be observed that 1) Regarding different model configurations, LSTM and TCN models will directly improve the model performance when increasing the model size, while only the models conditioned on TTFiLM and TVFiLM layers will improve in GCN and S4 models. We speculate this is because when these models are conditioned on the FiLM layer, the conditioning layer is not strong enough and will perform worse when increasing the model sizes. In contrast, the TFiLM layer is effective but has too many parameters, leading to learning problems when scaling up the model sizes. 2) Regarding different conditional layers, LSTM models with TVConcat perform significantly better than pseudo concatenation. In TCN, GCN, and S4 models, TVFiLM surprisingly performs the best, indicating the effectiveness of time-varying modeling when simulating analog compressors. It is also worth noting that the models conditioned with TFiLM generally perform second, followed by the TTFiLM models with similar performances, indicating its effectiveness in reducing model parameters and computational costs. 3) Regarding different model types, GCN generally performs the best, indicating the effectiveness of WaveNet-style dilated convolutions, followed by the S4 and TCN models conditioned on TTFiLM and TVFiLM. It is worth noting that the LSTM model with TVConcat outperforms many other baselines, showing the importance of the conditioning layer. 4) Regarding different test scenarios, LSTM models with TVConcat and TCN, GCN, and S4 models with TFiLM, TTFiLM, and TVFiLM achieve outstanding performance in both seen and unseen parameter settings. In contrast, a performance gap can be observed in simpler conditioning layers. This indicates the advanced generalization ability brought by the adequate temporal-varying conditional layers.
@@ 4.3 Grey-Box Methods
The benchmarking results on grey-box models are illustrated in Table 3. For pseudo-grey-box modeling, it can be observed that 1) Regarding different gain computer models, static gain with a soft knee generally performs better with different level detectors. This illustrates the model’s explainability since the analog module in the SSL-G-Bus compressor uses a soft knee where the knee width will automatically be computed by an internal algorithm based on the threshold and ratio  [^14] https://www.solidstatelogic.com/assets. 2) Regarding various level detectors, using the switching one-pole filter generally gives the best results, followed by using a simple one-pole filter. Using RNN-modulated one-pole filters, on the other hand, performs comparably worse. We speculate this is because, unlike the LA-2A compressor, where the optical circuit modules will bring a lot of non-linear distortion and coloration to the resulting audio, our VCA compressor uses simple level detectors; thus, switching to a more complex RNN-based model damages the performance. 3) Regarding different test sets, a comparable performance gap exists between the seen and unseen distributions. This is also because of the changing compressor curve in the analog module \@footnotemark, making it hard for grey-box models without feedback loops to learn that information. For analog simulation with effect chains, it is pretty surprising that using only two parametric gains for modeling the compression and two parametric EQs for modeling the non-linear distortions obtains the best performance — The fewer constraints give more improvement room for the non-linear deep learning modules. Experiments also show adding a simple phase inversion module would damage the model performance since there are no phasers in the actual analog module, confirming its effectiveness and explainability.
@@ 4.4 White-Box Plugins
To evaluate the development of academic NN-based models and further illustrate the effectiveness of our proposed dataset, benchmarking results on available white-box plugins are also reported, as shown in Table 4. It is worth noting that we do not conduct a comparative analysis between these plugins since they all come from different companies and are modeled in various environments, including using different analog devices from different manufacturers and using different audio interfaces, cables, and other hardware devices for optimization This will result in noticeable errors in evaluation, making the objective metric scores larger in multiple amounts. However, comparing these commercial plugins with existing academic black-box and grey-box models, a significant performance gap can still be observed, especially in extreme compression scenarios. Specifically, the plugin that comes from the PSPaudioware has an M-STFT score of 0.3047 and 0.2184 on the seen and unseen test sets, significantly outperforming the scores of 0.3238 and 0.3983 from the best academic NN-based models. This shows that the SOTA academic NN-based model in VA modeling still has a long way to go compared with industry plugins, which also illustrates the importance of our work since both model structure and datasets need to be improved for better performance.
@@ 4.5 Ablation Study
We also conducted ablation studies to illustrate the effectiveness of improving data quantity and diversity. We selected the GCN model conditioned with the TVFiLM layer as the baseline model and compared its performance when trained on different subsets. In particular, to control the data quantity, we fixed the number of total songs to 100 and control the length used to clip each song, resulting in 5 subsets from 3 minutes to 500 hours; to investigate the data diversity, we fixed the total data quantity to 50 hours and control the number of total songs with the adjusted clip lengths, resulting in 5 subsets from 5 songs to 100 songs. The detailed division and evaluation results are illustrated in Table 5. It can be observed that 1) increasing the data quantity steadily improves the model performance from 3 minutes to 500 hours, with the 50 hours as the division line for significant improvement, which is also confirmed by previous works [52]. 2) Increasing the data diversity is effective when there are only a few songs, and the improvement will be saturated until there are 50 different songs, especially in the unseen parameter settings. This is because different songs have different dynamics. Increasing the data diversity can help the model see more compression patterns, thus enhancing its generalization ability in unseen songs and parameter combinations.
@@ 5 Conclusion
In conclusion, this paper presents Diff-SSL-G-Comp, the first extensive and diverse dataset for DRC VA modeling. Our dataset comprises 2528 hours of processed unmastered songs in 220 parameter combinations with diverse genres, instruments, tempos, and keys. We provide benchmarking results on various open-sourced black-box and grey-box models, as well as available white-box plugins to facilitate the use of our dataset. We also provide ablation experiment results on different data subsets to illustrate the effectiveness of the improved data scale and quantity.
@@ 6 Acknowledgment
Table 5: Benchmarking results of existing commercial plugins on seen and unseen test sets. The best and second best results of every column are bold and underlined.
System L1 ($\downarrow$) M-STFT ($\downarrow$)
Seen Unseen Seen Unseen
Solid State Logic 0.0322 0.0175 0.4489 0.2943
Softube 0.0448 0.0237 0.7069 0.4546
Overloud 0.0326 0.0176 0.4738 0.3253
PSPaudioware 0.0269 0.0145 0.3047 0.2184
We acknowledge the computational resources provided by the Aalto Science-IT project. We acknowledge the EuroHPC Joint Undertaking for awarding this project access to the EuroHPC supercomputer LUMI, hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Regular Access call.

### v2

@@ 3.2 Black-Box Methods
Table 2 illustrates the benchmarking results on black-box methods. Several key observations can be made: 1) Regarding the effectiveness of parameter scaling, LSTM and TCN models consistently benefit from increased model size. In contrast, GCN and S4 models only improve when conditioned on TTFiLM or TVFiLM layers. We speculate that the baseline FiLM layers used in these models are not expressive enough, leading to degraded performance as model capacity increases. On the other hand, TFiLM is powerful but introduces too many parameters, which may cause training instability in larger models. 2) Regarding parameter efficiency across different models, the LSTM model with TVConcat at 8.0K parameters achieves competitive results compared to larger models, and the S4 models with TVFiLM and TTFiLM reach near SOTA performance under 12K parameters. In contrast, models using TFiLM layers often require significantly more parameters to achieve comparable performance, making them unsuited for resource-constrained environments. 3) Regarding different conditioning layers, LSTM models with TVConcat perform significantly better than with simple concatenation. For TCN, GCN, and S4 models, TVFiLM surprisingly achieves the best performance, highlighting the effectiveness of time-varying modulation in modeling analog compressors. TFiLM generally ranks second, followed closely by TTFiLM, which offers a favorable trade-off between performance and parameter efficiency. 4) Regarding different model types, GCN consistently outperforms other architectures, demonstrating the strength of WaveNet-style dilated convolutions. S4 and TCN models with TTFiLM or TVFiLM also perform well. Notably, LSTM models with TVConcat outperform many other baselines, emphasizing the importance of temporal conditioning. 5) Regarding the generalization ability to unseen test scenarios, LSTM models with TVConcat and TCN, GCN, and S4 models with TFiLM, TTFiLM, and TVFiLM maintain strong performance on both seen and unseen parameter settings. In contrast, models using simpler conditioning layers exhibit noticeable performance drops under unseen testing senarios.
@@ 3.3 Grey-Box Methods
The benchmarking results on grey-box models are presented in Table 3. It can be observed that: 1) Regarding different gain computer models, static gain with a soft knee generally performs better with different level detectors. This aligns with the analog design of the SSL G-Bus compressor, which employs a soft knee where the knee width is automatically computed based on the threshold and ratio  [^17] https://www.solidstatelogic.com/assets. 2) For different level detector implementations, the switching one-pole filter achieves the best overall performance, followed by the standard one-pole filter. In contrast, the RNN-modulated one-pole filter performs worse. We speculate that this is due to the relatively simple design of the VCA compressor’s level detection circuit, which is different from the LA-2A that has strong non-linear distortion due to its optical components. Under this scenario, overly complex models like RNN-based detectors may overfit and lead to quality degradation. 3) egarding different test sets, a noticeable performance gap is observed between seen and unseen parameter settings. This can also be attributed to the changing compressor curve in the analog module \@footnotemark, making it hard for grey-box models without explicit feedback mechanisms to capture that information. 4) In analog effect chain simulation, it is notable that the best performance is achieved using two parametric gain modules for compression and two parametric EQs for non-linear coloration. This illustrates the powerful learning ability of neural networks in loose conditions. Experiments also show adding a simple phase inversion module would damage the model performance since there are no phasers in the actual analog module, confirming its effectiveness and explainability.
@@ 3.4 White-Box Plugins
To evaluate the development of NN-based models and further illustrate the effectiveness of our proposed dataset, benchmarking results on white-box plugins are also reported, as shown in Table 4. Compared to these industry-standard plugins, a significant performance gap remains, particularly under extreme compression scenarios. This highlights that even the SOTA academic NN-based models still lag behind their commercial counterparts, which also illustrates the importance of our work since both model structure and datasets need to be improved for better performance.
@@ 3.5 Ablation Study
We also conducted ablation studies to illustrate the effectiveness of improving data quantity and diversity. We selected the GCN model conditioned with the TVFiLM layer as the baseline model and compared its performance when trained on different subsets. In particular, to control the data quantity, we fixed the number of total songs to 100 and control the length used to clip each song, resulting in 5 subsets from 3 minutes to 500 hours; to investigate the data diversity, we fixed the total data quantity to 50 hours and control the number of total songs with the adjusted clip lengths, resulting in 5 subsets from 5 songs to 100 songs. The detailed results are illustrated in Table 5. It can be observed that 1) increasing the data quantity steadily improves the model performance from 3 minutes to 500 hours, with the 50 hours as the division line for significant improvement, which is also confirmed by previous works [47]. 2) Increasing the data diversity is effective when there are only a few songs, and the improvement will be saturated until there are 50 different songs, especially in the unseen parameter settings.
@@ 4 Conclusion
In conclusion, this paper presents Solid State Bus-Comp, the first extensive and diverse dataset for DRC VA modeling. Our dataset comprises 2528 hours of processed unmastered songs in 220 parameter combinations with diverse genres, instruments, tempos, and keys. We provide benchmarking results on various open-sourced black-box and grey-box models, as well as available white-box plugins to facilitate the use of our dataset. We also provide ablation experiment results on different data subsets to illustrate the effectiveness of the improved data scale and quantity.
@@ 5 Acknowledgment
We acknowledge the computational resources provided by the Aalto Science-IT project. We acknowledge the EuroHPC Joint Undertaking for awarding this project access to the EuroHPC supercomputer LUMI, hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Regular Access call. This work is also supported by the 2023 Shenzhen stability Science Program and the Shenzhen Science and Technology Program (ZDSYS20230626091302006)